A BASKET CONTAINS 5 WHITE BALL (W) AND 5 BLACK BALL (B). WHAT IS THE POSSIBLE RESULT IF YOU DRAW 3 BALLS IN RANDOM FROM THE BASKET. LET X BE THE RANDOM VARIABLE REPRESENTING THE NUMBER OF BALCK BALL THAT OCCUR.





1
Expert's answer
2022-02-17T12:37:24-0500

"P(X=0)=\frac{5}{10}*\frac{4}{9}*\frac{3}{8}=\frac{1}{12}."

"P(X=1)=\frac{5}{10}*\frac{5}{9}*\frac{4}{8}*C_3^1=\frac{5}{12}."

"P(X=2)=\frac{5}{10}*\frac{4}{9}*\frac{5}{8}*C_3^2=\frac{5}{12}."

"P(X=3)=\frac{5}{10}*\frac{4}{9}*\frac{3}{8}=\frac{1}{12}."